Fission of one nucleus uranium-235 yields an average of approximately 200 MeV. What mass of uranium produces the energy of 3.30 x 10^10 J in fission?
1
Expert's answer
2021-02-22T10:29:05-0500

If one nucleus yields "200MeV=3.2\cdot10^{-11}J", then we need "N=\frac{3.3\cdot 10^{10}}{3.2\cdot10^{-11}}\approx 1.03\cdot10^{21}" of nuclei. The molar mass of U-235 is "235 g\/mol", where "1 mol" contains "6.02\cdot10^{23}" atoms/molecules. Therefore the mass of "N" U-235 nuclei (the mass of a nucleus is approximately the mass of an atom, as the electron mass is neglected) is "235\cdot\frac{1.03\cdot10^{21}}{6.02\cdot10^{23}}\approx 0.4 g".